X is a normally distributed random variable with mean equal to 20 and variance equal to 100. The probability that X is < 30 is equal to the probability that Z is less than:
Answer #1

\mu = Mean = 20

\sigma = SD = \sqrt{100}=10

P(X<30):
Transforming to Standard Normal Variate:

we get:

Z = (X - \mu )/\sigma

= (30 -20)/10 = 1

So,

Answer is:

1